You are considering two investment options. In option A, you have to invest RM5000 now and RM1000 three years from now, In option B, you have to invest RM2500 now, RM1500 a year from now, and RM1000 three years from now. In both options, you will receive four annual payments of RM3000 each. (You will get the first payment a year from now.) Which of these options would you choose based on (a) the conventional payback criterion, and (b) the present worth criterion, assuming 7%) interest? Based on conventional payback period method, choose either Format: A (A/B/Both) Format PW Option A 8439.9 Format PW Option B 9986.8 Based on PW analysis method, choose Option (A or B) Format A
Answer #1

Answer

Option A

Initial Investment = $5000

Investment at 3rd year = $1000

Annual Benefits = $3000 each for 4 years

Rate of interest is 7%

NPW = -$5000 – $1000 (P/F, 7%, 3) + $2000 (P/A, 7%, 4)

NPW = -$5000 – $1000 (0.81630) + $2000 (0.76290)

= 2657.9

AE = NPW (A/P, 7%, 4)

AE = 2657.9 (A/P, 7%, 4)

AE = 1088.49 (0.76290)

AE = 830.40

Option B

Initial Investment = $2500

Investment at 1st year = $1500

Investment at 3rd year = $1000

Annual Benefits = $3000 each for 4 years

Rate of interest is 7%

NPW = -$2500 – $1500 (P/F, 7%, 1) – $1000 (P/F, 7%, 3) + $3000 (P/A, 7%, 4)

NPW = -$2500 – $1500 (0.93458) – $1000 (0.81630) + $3000 (0.76290)

= 2570.53

AE = NPW (A/P, 7%, 4)

AE = 2570.53 (A/P, 7%, 4)

AE = 2570.53 (0.76290)

AE = 1961.057337

OPTION B is a better option as it has AE more than option A.